OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1

OpenStudy (alexwee123):

distribute out the 7 and 5 then isolate x

OpenStudy (anonymous):

7(x − 1) = 5(x + 2) + 1 7x - 7 = 5x + 10 + 1 7x - 7 = 5x + 11 2x = 18 x=9 To test that this is correct, plug x back into the equation.. 7(9-1) = 5(9+2) + 1? 7(8) = 5(11) + 1 56 = 56

OpenStudy (anonymous):

1) distribute each of the numbers, so the eqaution becomes 7x-7=5x+10+1 2) Notice how in the second part you can combine like terms to make it 7x-7=5x+11 3) Then subtract 5x from both sides to get 2x-7=11 4) Then add 7 to both sides to get 2x=18 5) Just divide both sides by 2 to get x=9 hope that helps :)